Determine the required value of the missing probability to make the distribution a discrete probability distribution

x

​P(x)   

3

0.34

4

​?

5

0.33

6

0.16

P(4)=
Answer #1

solution:.

Solution :

using properties of probability distribution

x

​P(x)   

3

0.34

4

​?

5

0.33

6

0.16

P(4)=?

P(X=x)  = P(X=x) = 1

=p(x=3)+p(x=4)+p(x=5)+p(x=6)=1

=0.34+p(x=4)+0.33+0.16=1

=0.83+p(x=4)=1

=p(x=4)=1-0.83=0.17

=p(x=4)=0.17

p(4)=0.14

=